OpenStudy (spareb665):

Solve the equation 2^3x+1=5^-x and round your answer to two digits after the decimal sign.

OpenStudy (spareb665):



OpenStudy (p0sitr0n):

add log on both sides

OpenStudy (p0sitr0n):

(3x+1)log2 = -xlog5 3xlog2 +log2 = -xlog5 3xlog2 + xlog5 = -log2 x(3log2+log5) = -log2 x = -log2/(3log2+log5)

OpenStudy (phi):

use the idea that \[ \log(a^b) = b\ \log(a) \] you can use any log, but log base 10 or natural log ln are available on a calculator